Please help! I am having a hard time figuring out this question:
The demand equation for a certain type of printer is given by D = -200p +35,000. The supply equation is predicted to be S = -p^2 +400p - 20,000. What is the equilibrium price?
Thank you!!

Answer by rapaljer(4671)

Equilibrium is attained when supply equals demand. Therefore
D = -200p +35,000
S = -p^2 +400p - 20,000
-200p +35,000 = -p^2 +400p - 20,000

Set the quadratic equation equal to zero, and hope it factors! If it does not, then you will have to use the quadratic formula or completing the square (or a graphing calculator!) to get the answer.

p^2 -600p+ 55000 = 0

This does NOT factor, so the quadratic formula, completing the square, or graphing calculator will be needed. The answers I got with calculator (rounded to nearest whole number) are p=113 or p=487.

Life doesn't always come out even. Nevertheless, the answers in math so often are "rigged" to work out that I am suspicious when they do NOT come out even. So, check my work, and see if I am correct!! Let me know if you need help solving the quadratic equation.

R^2 at SCC